OpenStudy (kiakat):

Help! Simplify: (sin Θ − cos Θ)^2 − (sin Θ - cos Θ)^2

OpenStudy (anonymous):

sin2Θ

OpenStudy (kiakat):

how'd you get that though...

OpenStudy (triciaal):

sin^2 -2sin cos + cos^2 -(sin^2 -2sin Cos + cos^2) sin^2 -2sincos + cos^2 -sin^2 +2sinCos -cos^2

OpenStudy (anonymous):

2*sin(Θ)*cos(Θ) = sin(2Θ) sin(Θ)*cos(Θ) = (1/2)*sin(2Θ)

Nnesha (nnesha):

you don't have to distribute but good to do it for practice both parentheses are same so let's say (sin Θ − cos Θ)^2 = x^2 so \[\huge\rm x^2 -x^2 = ?\]

OpenStudy (kiakat):

so the answer would be zero?

Nnesha (nnesha):

yes right

OpenStudy (kiakat):

okay, thank you guys!!!

Nnesha (nnesha):

np